# Preferences Evolve And So Should Your Bandits: Bandits with Evolving States for Online Platforms

### v2

@@ 15             Play arm $i$ Play arm $j$
Fix two arms $i,j\in[K]$ with $i\neq j$, $\lambda>0$ and a scalar $\varepsilon>0$. Assume that at some round $s$, after a history of play $H^{\prime}$, we are at state $q_{s}$. Then, playing alternately arms $i,j$ for infinitely many rounds $t$ makes the state become:
[…]
@@
Let $\textrm{ALT}_{i,j}$ be the algorithm that continuously alternates between arm $i$ and $j$. We first prove that :
$q_{s+t+1}\left(H^{ALT_{i,j}}_{s:s+t}\right)=\begin{cases}(1-\lambda)^{t}q_{s+1}+\lambda\sum_{\tau=0}^{(t-2)/2}(1-\lambda)^{2\tau}b_{j}+(1-\lambda)^{2\tau+1}b_{i}&\textrm{if }t\equiv 0\pmod{2},\\ (1-\lambda)^{t}q_{s+1}+\lambda\sum_{\tau=0}^{(t-3)/2}(1-\lambda)^{2\tau+1}b_{j}+\sum_{\tau=0}^{(t-1)/2}(1-\lambda)^{2\tau}b_{i}&\textrm{else}\end{cases},$ (30)
where $t\geq 0$, using induction. For the base case rounds $s+1$, $s+2$ the state becomes:
$q_{s+1}\left(H^{ALT_{i,j}}_{s:s}\right)=q_{s+1}\quad\textrm{and}\quad q_{s+2}\left(H^{ALT_{i,j}}_{s:s+1}\right)=(1-\lambda)q_{s+1}+\lambda b_{i}.$
@@
$q_{s+n+1}\left(H^{ALT_{i,j}}_{s:s+n}\right)=(1-\lambda)^{n}q_{s+1}+\lambda\sum_{\tau=0}^{(n-2)/2}(1-\lambda)^{2\tau}b_{j}+(1-\lambda)^{2\tau+1}b_{i},$
@@
$q_{s+n+2}\left(H^{ALT_{i,j}}_{s:s+n+1}\right)$ $=(1-\lambda)q_{s+n+1}+\lambda b_{i}$
@@
$=(1-\lambda)^{t}+\lambda\sum_{\tau=0}^{(t-2)/2}(1-\lambda)^{2\tau}b_{j}+(1-\lambda)^{2\tau+1}b_{j}$
@@
Putting in Eq. (30) we get:
$q_{s+t+1}\left(H^{ALT_{i,j}}_{s:s+t}\right)=\begin{cases}(1-\lambda)^{t}q_{s+1}+\frac{1-(1-\lambda)^{t}}{2-\lambda}b_{j}+(1-\lambda)\frac{1-(1-\lambda)^{t}}{2-\lambda}b_{i}&\textrm{if }t\equiv 0\pmod{2},\\ (1-\lambda)^{t}q_{s+1}+(1-\lambda)\frac{1-(1-\lambda)^{t-2}}{2-\lambda}b_{j}+\frac{1-(1-\lambda)^{t-2}}{2-\lambda}b_{i}&\textrm{else}\end{cases},$
It is obvious that take the limit to infinity in Eq. 30:
$\lim_{t\to\infty}q_{s+t}=\begin{cases}b_{i}\frac{1-\lambda}{2-\lambda}+b_{j}\frac{1}{2-\lambda}&\textrm{if}\ t\equiv 0\pmod{2}\\ b_{j}\frac{1-\lambda}{2-\lambda}+b_{i}\frac{1}{2-\lambda}&\textrm{else}\\ \end{cases}.$
Using the Lemma 3.9 we can play $N(\lambda)$ rounds two arms $i,j$ alternately and get state:
$\left|q_{s+2N(\lambda)}-b_{i}\frac{1-\lambda}{2-\lambda}-b_{j}\frac{1}{2-\lambda}\right|\leq\varepsilon.\quad\textrm{if}\ t\equiv 0\pmod{2}$
$\left|q_{s+2N(\lambda)}-b_{j}\frac{1-\lambda}{2-\lambda}-b_{i}\frac{1}{2-\lambda}\right|\leq\varepsilon.\quad\textrm{if}\ t\equiv 1\pmod{2}$
@@
For each $N(\lambda)$ the estimate $\hat{\lambda}$ is between:
$1+\frac{\mathbb{E}[\hat{r}_{i,b_{j}}]-\mathbb{E}[\hat{r}_{i,j}]}{\mathbb{E}[\hat{r}_{i,b_{i}}]-\mathbb{E}[\hat{r}_{i,j}]}-\Omega(\delta)\leq\hat{\lambda}\leq 1+\frac{\mathbb{E}[\hat{r}_{i,b_{j}}]-\mathbb{E}[\hat{r}_{i,j}]}{\mathbb{E}[\hat{r}_{i,b_{i}}]-\mathbb{E}[\hat{r}_{i,j}]}+O(\delta),$
with probability at least $4\exp(-2M\delta^{2})$ under the assumption of $(A):|\mathbb{E}[\hat{r}_{i,b_{j}}]-\mathbb{E}[\hat{r}_{i,j}]|\geq o(\delta)$ and $|\mathbb{E}[\hat{r}_{i,i}]-\mathbb{E}[\hat{r}_{i,j}]|-2\delta\geq o(\delta)$. Where $\hat{r}_{i,i}$ is the estimator that algorithm builds of random variable $R_{i,i}\sim\textsf{Bern}\left(r_{i}\cdot\left((1-\lambda)^{\tilde{N}(\lambda)}(q_{0,i}-b_{i})+b_{i}\right)\right)$, $q_{0,i}$ is the value of state before we start sampling. Respectively, $\hat{r}_{i,b_{j}}$ is the estimator that algorithm builds of random variable $R_{i,b_{j}}\sim\textsf{Bern}\left(r_{i}\cdot\left((1-\lambda)^{\tilde{N}(\lambda)}(q_{0,i}-b_{j})+b_{j}\right)\right)$, $q_{0,b_{j}}$ is the value of state before we start sampling. And $\hat{r}_{i,j}$ is the estimator that algorithm builds of random variable $R_{i,j}\sim\textsf{Bern}\left(r_{i}\cdot\left((1-\lambda)^{\tilde{N}(\lambda)}(q_{0,j}-\frac{b_{j}+(1-\lambda)b_{i}}{2-\lambda})+\frac{b_{j}+(1-\lambda)b_{i}}{2-\lambda}\right)\right)$ where $q_{0,j}$ is the value of state before we start sampling.
@@
For the estimator $\hat{r}_{i,b_{j}}$ we begin from a random state $q_{0,b_{j}}$ and then the we play arm $j$ for $\tilde{N}(\lambda)$ rounds so the state becomes from Lemma 3.5:
$q_{t}=(1-\lambda)^{\tilde{N}(\lambda)}(q_{0,b_{j}}-b_{j})+b_{j}.$
Using Hoeffding’s Inequality as in Lemma 3.8 we get:
$\Pr\left[\left|\hat{r}_{i,b_{j}}-r_{i}\cdot\left((1-\lambda)^{\tilde{N}(\lambda)}(q_{0,b_{j}}-b_{j})+b_{j}\right)\right|\geq\delta\right]\leq 2\exp\left(-2M\delta^{2}\right)$
Then, for $\hat{r}_{i,i}$ we fist play arm $j$ for $\tilde{N}(\lambda)$ rounds and the state equals to another random value $q_{0,i}$. Then, we play arm $i$ for $\tilde{N}(\lambda)$ rounds and the state becomes $q_{t}=(1-\lambda)^{\tilde{N}(\lambda)}(q_{0,i}-b_{i})+b_{i}$ and then we take a sample from $\textsf{Bern}(r_{i}q_{t})$. Thus, the expected value of $\mathbb{E}[\hat{r}_{i,i}]=r_{i}\cdot\left((1-\lambda)^{\tilde{N}(\lambda)}(q_{r}-b_{i})+b_{i}\right)$, and using Hoeffding’s inequality we get:
$\Pr\left[\left|\hat{r}_{i,j}-r_{i}\cdot\left((1-\lambda)^{\tilde{N}(\lambda)}(q_{0,i}-b_{i})+b_{j}\right)\right|\geq\delta\right]\leq 2\exp\left(-2M\delta^{2}\right)$
Then the same applies for $\hat{r}_{i,j}$:
$\Pr\left[\left|\hat{r}_{i,j}-r_{i}\cdot\left((1-\lambda)^{\tilde{N}(\lambda)}\left(q_{0,j}-\frac{b_{j}+(1-\lambda)b_{i}}{2-\lambda}\right)+\frac{b_{j}+(1-\lambda)b_{i}}{2-\lambda}\right)\right|\geq\delta\right]\leq 2\exp\left(-2M\delta^{2}\right)$
Thus, $\hat{\lambda}$ is upper bounded by:
$\hat{\lambda}$ $=\frac{\hat{r}_{i,b_{i}}+\hat{r}_{i,b_{j}}-2\hat{r}_{i,j}}{\hat{r}_{i,b_{i}}-\hat{r}_{i,j}}$
$=1-\left|\frac{\hat{r}_{i,b_{j}}-\hat{r}_{i,j}}{\hat{r}_{i,b_{i}}-\hat{r}_{i,j}}\right|\leq 1-\frac{|\mathbb{E}[\hat{r}_{i,b_{j}}]-\mathbb{E}[\hat{r}_{i,j}]|-2\delta}{|\mathbb{E}[\hat{r}_{i,b_{i}}]-\mathbb{E}[\hat{r}_{i,j}]|+2\delta}$ ( $\frac{\hat{r}_{i,b_{j}}-\hat{r}_{i,j}}{\hat{r}_{i,b_{i}}-\hat{r}_{i,j}}\leq 0$ as ($\lambda\leq 1$))
$\leq 1-\frac{|\mathbb{E}[\hat{r}_{i,b_{j}}]-\mathbb{E}[\hat{r}_{i,j}]|}{|\mathbb{E}[\hat{r}_{i,b_{i}}]-\mathbb{E}[\hat{r}_{i,j}]|}+O(\delta)$ (From assumption (A))
$=1+\frac{\mathbb{E}[\hat{r}_{i,b_{j}}]-\mathbb{E}[\hat{r}_{i,j}]}{\mathbb{E}[\hat{r}_{i,b_{i}}]-\mathbb{E}[\hat{r}_{i,j}]}+O(\delta)$
and lower bounded by
$\hat{\lambda}$ $=\frac{\hat{r}_{i,b_{i}}+\hat{r}_{i,b_{j}}-2\hat{r}_{i,j}}{\hat{r}_{i,b_{i}}-\hat{r}_{i,j}}$
$=1-\left|\frac{\hat{r}_{i,b_{j}}-\hat{r}_{i,j}}{\hat{r}_{i,b_{i}}-\hat{r}_{i,j}}\right|\geq 1-\frac{|\mathbb{E}[\hat{r}_{i,b_{j}}]-\mathbb{E}[\hat{r}_{i,j}]|+2\delta}{|\mathbb{E}[\hat{r}_{i,b_{i}}]-\mathbb{E}[\hat{r}_{i,j}]|-2\delta}$ ( $\frac{\hat{r}_{i,b_{j}}-\hat{r}_{i,j}}{\hat{r}_{i,b_{i}}-\hat{r}_{i,j}}\leq 0$ as ($\lambda\leq 1$))
$=1-\frac{|\mathbb{E}[\hat{r}_{i,b_{j}}]-\mathbb{E}[\hat{r}_{i,j}]|}{|\mathbb{E}[\hat{r}_{i,b_{i}}]-\mathbb{E}[\hat{r}_{i,j}]|}-\Omega(\delta)$ (From assumption $(A)$)
$=1+\frac{\mathbb{E}[\hat{r}_{i,b_{j}}]-\mathbb{E}[\hat{r}_{i,j}]}{\mathbb{E}[\hat{r}_{i,b_{i}}]-\mathbb{E}[\hat{r}_{i,j}]}-\Omega(\delta)$
These are satisfied with probability at least $4\exp(-2M\delta)$. ∎
Now we move on the main Algorithm 8. First, we take some samples to distinguish if $\lambda$ is near 00 or not and if the $r_{i}$’s and the gaps $b_{i}-b_{j}$ are small. If it is near 00 we call $EXP3.P$ otherwise, we try $\tilde{N}(\lambda)$s in the Algorithm 7 until we find the right one.
/* Separate to small and big $\lambda$/
1 for each arm $i\in[K]$ do
       Play arm $i$ for $(T/K)^{2/3}$ rounds. // arm $i$ chosen at random
       /* State now is at: $q_{t}\approx b_{i}$, if $\lambda$ is big enough */
2       for $(T/K)^{2/3}$ rounds do
             Play arm $i$ (same as in Line 1) and update estimator $\hat{\mu}_{i}\leftarrow\hat{r}_{i}+R_{i}/(T/K)^{2/3}$. // $rew_{t}\sim\textsf{Bern}(q_{t}r_{i})$
3
4Choose a random arm $j\in[K]$. for each $i\in[K]\setminus j$ do
5       Play arms $i,j$ alternately for $2(T/K)^{2/3}$ rounds. for $2(T/K)^{2/3}$ rounds do
6             Play arm $i$ and update $\hat{\mu}_{(i,j)}\leftarrow\hat{\mu}_{(i,j)}+R_{i,j}/\sqrt{T}$. Play arm $j$.
/* Small $\lambda$. */
7 if $|\hat{\mu}_{i}-\hat{\mu}_{i,j}|\leq 3\frac{\sqrt{\log T}}{(T/K)^{1/3}}$ then
8       Call EXP3.P
/* Big $\lambda$. */
9 else
10       $\tilde{N}(\lambda)\leftarrow\log T$ Call Algorithm 7 for $\tilde{N}(\lambda)$, arms $i,j$ and get $\hat{\lambda}_{1}$ Call Algorithm 7 for $2\tilde{N}(\lambda)$, arms $i,j$ and get $\hat{\lambda}_{2}$ $\tilde{N}(\lambda)\leftarrow 4\tilde{N}(\lambda)$ while $|\hat{\lambda}_{1}-\hat{\lambda}_{2}|\geq\delta$ do
11             Call Algorithm 7 for $\tilde{N}(\lambda)$, arms $i,j$ and get $\hat{\lambda}_{3}$ $\hat{\lambda}_{1}\leftarrow\hat{\lambda}_{2}$ $\hat{\lambda}_{2}\leftarrow\hat{\lambda}_{3}$ $\tilde{N}(\lambda)\leftarrow 2\tilde{N}(\lambda)$
12      Call Algorithm 2 for $\lambda=\hat{\lambda}_{1}$
ALGORITHM 8 B-DES Unknown $\lambda$
In lemmas E.5 and lemmas E.6 we bound the estimators $\hat{\mu}_{i}$ and $\hat{\mu}_{i,j}$. Then, we bound their difference and prove that if real $\lambda$ is small we call $EXP3.P$ otherwise we are trying to learn real $\lambda$.
@@ 12      Call Algorithm 2 for $\lambda=\hat{\lambda}_{1}$
For the estimators $\hat{\mu}_{i}$ it holds:
$\Pr\left[|\hat{\mu}_{i}-Y_{i}|\geq\frac{\sqrt{\log T}}{(T/K)^{1/3}}\right]\leq\frac{1}{T^{2}},$
@@
$Y_{i}=r_{i}\left((1-\lambda)^{(T/K)^{2/3}}\frac{1-(1-\lambda)^{(T/K)^{2/3}}}{T^{2/3}\lambda}(q_{0,i}-b_{i})+b_{i}\right)$
and $q_{0,i}$ the value of state before we start building the estimator.
@@
Let $q_{0,i}$ be the state before we start pulling arm $i$. Then, after $(T/K)^{2/3}$ pulls of this arm the state becomes:
$q_{(T/K)^{2/3}+1,i}$ $=(1-\lambda)^{(T/K)^{2/3}}+\lambda\sum_{s=0}^{T^{2/3}-1}(1-\lambda)^{s}b_{i}$
$=(1-\lambda)^{(T/K)^{2/3}}+\lambda\cdot\frac{1-(1-\lambda)^{(T/K)^{2/3}}}{1-(1-\lambda)}b_{i}$
$=(1-\lambda)^{(T/K)^{2/3}}(q_{0,i}-b_{i})+b_{i}$
Then, for another $(T/K)^{2/3}$ rounds the algorithm samples reward from $\textsf{Bern}(r_{i}\cdot q_{(T/K)^{2/3}+s,i})$ for $s\in[(T/K)^{2/3}]$, where for each $s\in[(T/K)^{2/3}]$ it applies:
$q_{(T/K)^{2/3}+s,i}$ $=(1-\lambda)^{(T/K)^{2/3}+s}\cdot(q_{s,i}-b_{i})+b_{i}$
Thus, the expected reward for each $s\in[(T/K)^{2/3}]$ is $\mathbb{E}[R_{i,s}]=r_{i}\cdot\left((1-\lambda)^{(T/K)^{2/3}+s}\cdot(q_{0,i}-b_{i})+b_{i}\right)$, and
$\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}q_{t}\right]$ $=\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}\cdot\left((1-\lambda)^{(T/K)^{2/3}+s}\cdot(q_{0,i}-b_{i})+b_{i}\right)$
$=r_{i}\left((1-\lambda)^{(T/K)^{2/3}}\frac{1-(1-\lambda)^{(T/K)^{2/3}}}{\lambda}(q_{0,i}-b_{i})+(T/K)^{2/3}b_{i}\right)$
Using Hoeffding’s inequality on $\hat{\mu}_{i}$ we get the result.
@@
For the estimators $\hat{\mu}_{i,j}$ it applies:
$\Pr\left[|\hat{\mu}_{i,j}-Y_{i,j}|\geq\frac{\sqrt{\log T}}{(T/K)^{1/3}}\right]\leq\frac{1}{T^{2}},$
where
$Y_{i,j}=\frac{(1-\lambda)^{2(T/K)^{2/3}}}{(T/K)^{2/3}}\left(r_{i}q_{0,(i,j)}-\frac{r_{i}b_{j}}{2-\lambda}-(1-\lambda)\frac{r_{i}b_{i}}{2-\lambda}\right)\frac{1-(1-\lambda)^{2(T/K)^{2/3}}}{1-(1-\lambda)^{2}}+\frac{r_{i}b_{j}+(1-\lambda)r_{i}b_{i}}{2-\lambda},$
and $q_{0,(i,j)}$ the value of state before we start building the estimator.

### v4

@@ E.2.2 Analysis
Fix two arms $i,j\in[K]$ with $i\neq j$, $\lambda>0$ and a scalar $\varepsilon>0$. Assume that at some round $s$, after a history of plays $H^{\prime}_{1:s-1}$, we are at state $q_{s}$. Then, if $\lambda>0$, playing alternately arms $i,j$ for infinitely many rounds $t$ the state asymptotically becomes:
$q_{s+t+1}=\begin{cases}W_{i,j}(\lambda)&\textrm{if}\ t\equiv 0\pmod{2}\\ W_{j,i}(\lambda)&\textrm{otherwise}\\ \end{cases}.$
If $\lambda=0$, then the state is always $q_{s}=q_{0}$.
@@
Let $\textrm{ALT}_{i,j}$ be the algorithm that continuously alternates between arm $i$ and $j$. We first use induction to prove that:
$q_{s+t+1}\left(H^{\textrm{ALT}_{i,j}}_{s:s+t}\right)=\begin{cases}(1-\lambda)^{t}q_{s+1}+\lambda\sum_{\tau=0}^{(t-2)/2}\left[(1-\lambda)^{2\tau}b_{j}+(1-\lambda)^{2\tau+1}b_{i}\right],&\textrm{if }t\equiv 0\pmod{2},\\ (1-\lambda)^{t}q_{s+1}+\lambda\sum_{\tau=0}^{(t-3)/2}\left[(1-\lambda)^{2\tau+1}b_{j}+\sum_{\tau=0}^{(t-1)/2}(1-\lambda)^{2\tau}b_{i}\right],&\textrm{otherwise}\end{cases}$ (31)
where $t\geq 0$. For the base case rounds $s+1$, $s+2$ the state becomes:
$q_{s+1}\left(H^{\textrm{ALT}_{i,j}}_{s:s}\right)=q_{s+1}\quad\textrm{and}\quad q_{s+2}\left(H^{\textrm{ALT}_{i,j}}_{s:s+1}\right)=(1-\lambda)q_{s+1}+\lambda b_{i}.$
@@
$q_{s+n+1}\left(H^{\textrm{ALT}_{i,j}}_{s:s+n}\right)=(1-\lambda)^{n}q_{s+1}+\lambda\sum_{\tau=0}^{(n-2)/2}(1-\lambda)^{2\tau}b_{j}+(1-\lambda)^{2\tau+1}b_{i},$
@@
$q_{s+n+2}\left(H^{\textrm{ALT}_{i,j}}_{s:s+n+1}\right)$ $=(1-\lambda)q_{s+n+1}+\lambda b_{i}$
@@
$=(1-\lambda)^{t}+\lambda\sum_{\tau=0}^{(t-2)/2}(1-\lambda)^{2\tau}b_{j}+(1-\lambda)^{2\tau+1}b_{j}{}$ (32)
@@
Substituting the above in Eq. (32), we get:
$q_{s+t+1}\left(H^{\textrm{ALT}_{i,j}}_{s:s+t}\right)=\begin{cases}(1-\lambda)^{t}q_{s+1}+\frac{1-(1-\lambda)^{t}}{2-\lambda}b_{j}+(1-\lambda)\frac{1-(1-\lambda)^{t}}{2-\lambda}b_{i}&\textrm{if }t\equiv 0\pmod{2},\\ (1-\lambda)^{t}q_{s+1}+(1-\lambda)\frac{1-(1-\lambda)^{t-2}}{2-\lambda}b_{j}+\frac{1-(1-\lambda)^{t-2}}{2-\lambda}b_{i}&\textrm{else}\end{cases},$
Note that if $\lambda=0$, then $q_{s+t+1}\left(H^{\textrm{ALT}_{i,j}}_{s:s+t}\right)=q_{s+1}$, which is to be expected by the definition of our setting. Now, taking the limit $t\to\infty$ in Eq. (31), we get:
$\lim_{t\to\infty}q_{s+t}=\begin{cases}b_{i}\frac{1-\lambda}{2-\lambda}+b_{j}\frac{1}{2-\lambda}&\textrm{if}\ t\equiv 0\pmod{2}\\ b_{j}\frac{1-\lambda}{2-\lambda}+b_{i}\frac{1}{2-\lambda}&\textrm{otherwise}\\ \end{cases}.$
which concludes our proof. ∎
Combining Lemma E.3 with Lemma 3.10, we can alternate between arms $i,j$ for $N(\lambda)=\log(\varepsilon)/\log(1-\lambda)$ rounds and then with high probability we have that:
$\left|q_{s+N(\lambda)+t}-W_{i,j}(\lambda)\right|\leq\varepsilon,\quad\textrm{if}\ t\equiv 0\pmod{2}$
$\left|q_{s+N(\lambda)+t}-W_{i,j}(\lambda)\right|\leq\varepsilon,\quad\textrm{if}\ t\equiv 1\pmod{2}$
In Lemmas E.4 and E.5, we analyze how good the estimators $\hat{\mu}_{i}$ and $\hat{\mu}_{i,j}$ are. As a reminder, $\hat{\mu}_{i,j}$ is the estimator obtained by playing arm $j$ after arm $i$ has been played for $(T/K)^{2/3}$ rounds.
@@
For estimator $\hat{\mu}_{i}$ that Algorithm 8 builds, it holds that:
$\Pr\left[\left|\hat{\mu}_{i}-\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]\right|\leq\frac{\sqrt{\log T}}{(T/K)^{1/3}}\right]\geq 1-\frac{1}{T^{2}},$
where
$\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]=\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}\cdot\left((1-\lambda)^{(T/K)^{2/3}+s}\cdot(q_{0,i}-b_{i})+b_{i}\right),$
$R_{s,i}$ is the $s$-th reward sample from $\textsf{Bern}(r_{i}q_{t})$ when arm $i$ is pulled in line $2$ and $q_{0,i}$ is the value of starting state before starting pull arm $i$.
@@
Let $q_{0,i}$ be the value of the state before we start pulling arm $i$ for $(T/K)^{2/3}$ rounds. Then, after $(T/K)^{2/3}$ pulls of this arm the state becomes:
$q_{(T/K)^{2/3}+1,i}$ $=(1-\lambda)^{(T/K)^{2/3}}q_{0,i}+\lambda\sum_{s=0}^{T^{2/3}-1}(1-\lambda)^{s}b_{i}$
$=(1-\lambda)^{(T/K)^{2/3}}q_{0,i}+\lambda\cdot\frac{1-(1-\lambda)^{(T/K)^{2/3}}}{1-(1-\lambda)}b_{i}$
$=(1-\lambda)^{(T/K)^{2/3}}(q_{0,i}-b_{i})+b_{i}$
Then, for the next $(T/K)^{2/3}$ rounds the algorithm pulls arm $i$ and samples reward from
$R_{s,i}\sim\textsf{Bern}(r_{i}\cdot q_{(T/K)^{2/3}+s,i})$ for $s\in[(T/K)^{2/3}]$. For each $s\in[(T/K)^{2/3}]$, the state is:
$q_{(T/K)^{2/3}+s,i}$ $=(1-\lambda)^{(T/K)^{2/3}+s}\cdot(q_{0,i}-b_{i})+b_{i}.$
The expected reward collected for rounds at each $s\in[(T/K)^{2/3}]$ is
$\mathbb{E}[R_{s,i}]=\mathbb{E}\left[R_{s,i}\sim\textsf{Bern}(r_{i}\cdot q_{(T/K)^{2/3}+s,i})\right]$ $=r_{i}\cdot\left((1-\lambda)^{(T/K)^{2/3}+s}\cdot(q_{0,i}-b_{i})+b_{i}\right)$
Summing for all $s\in[K]$:
$\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]$ $=\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}\cdot\left((1-\lambda)^{(T/K)^{2/3}+s}\cdot(q_{0,i}-b_{i})+b_{i}\right)$
Using union bound over all arms we have Using Hoeffding’s inequality on $\hat{\mu}_{i}$ we get the following result:
$\Pr\left[\left|\hat{\mu}_{i}-\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]\right|\leq\frac{\sqrt{\log T}}{(T/K)^{1/3}}\right]\geq 1-\frac{1}{T^{2}},$ (33)
Using union bound over all arms we have the result. ∎
@@
Fix a pair $(i,j)$, where $i,j\in[K]$. Then, for the estimators $\hat{\mu}_{i,j}$ that Algorithm 8 builds it holds that:
$\Pr\left[\left|\hat{\mu}_{i,j}-\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{T^{2/3}-1}R_{s,i}^{\text{ALT}}\right]\right|\geq\frac{\sqrt{\log T}}{(T/K)^{1/3}}\right]\leq\frac{1}{T^{2}},$
@@
$\mathbb{E}\left[\sum_{s=0}^{T^{2/3}-1}R_{s,i}^{\text{ALT}}\right]$ $=\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}\cdot(1-\lambda)^{(T/K)^{2/3}+2s}\left(q_{0,i}^{\text{ALT}}-W_{i,j}(\lambda)\right)+W_{i,j}(\lambda),$
$q_{0,i}^{\text{ALT}}$ is the value of state before we start alternating arms and $R_{s,i}^{\text{ALT}}$ is the reward sampled during the alternation between arms $i,j$.
@@
Algorithm 8 first alternates between arm $i$ and arm $j$ for $(T/K)^{2/3}$ rounds. After this many rounds, the state becomes (see Lemma E.3):
$q_{(T/K)^{2/3}+1,i}^{\text{ALT}}$ $=(1-\lambda)^{(T/K)^{2/3}}q_{0,i}^{\text{ALT}}+\frac{1-(1-\lambda)^{(T/K)^{2/3}}}{2-\lambda}b_{j}+(1-\lambda)\frac{1-(1-\lambda)^{(T/K)^{2/3}}}{2-\lambda}b_{i}$
$=(1-\lambda)^{(T/K)^{2/3}}\left(q_{0,i}^{\text{ALT}}-W_{i,j}(\lambda)\right)+W_{i,j}(\lambda)$
Next, for another $2(T/K)^{2/3}$ rounds the algorithm keeps alternating between the two arms and samples reward from $R_{s,i}^{\text{ALT}}\sim\textsf{Bern}(r_{i}\cdot q_{(T/K)^{2/3}+2s+1,i}^{\text{ALT}})$, where
$q_{(T/K)^{2/3}+2s+1,i}^{\text{ALT}}$ $=(1-\lambda)^{(T/K)^{2/3}+2s}q_{0,i}^{\text{ALT}}+\frac{1-(1-\lambda)^{(T/K)^{2/3}+2s}}{2-\lambda}b_{j}+(1-\lambda)\frac{1-(1-\lambda)^{(T/K)^{2/3}+2s}}{2-\lambda}b_{i}$
$=(1-\lambda)^{(T/K)^{2/3}+2s}\left(q_{0,i}^{\text{ALT}}-W_{i,j}(\lambda)\right)+W_{i,j}(\lambda)$
Using this expression for the state, the expected reward obtained is equal:
$\mathbb{E}\left[\sum_{s=0}^{T^{2/3}-1}R_{s,i}^{\text{ALT}}\right]$ $=\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}\cdot(1-\lambda)^{(T/K)^{2/3}+2s}\left(q_{0,i}^{\text{ALT}}-W_{i,j}(\lambda)\right)+r_{i}W_{i,j}(\lambda)$
Finally, using Hoeffding’s inequality as in Lemma E.4 on $\hat{\mu}_{i,j}$, we get the result. ∎
Taking a union bound over all $i\in[K]$ and the estimators $\hat{\mu}_{i}$ and $\hat{\mu}_{i,j}$ they correspond to, we get that Lemmas E.4 and E.5 will hold for all $i,j\in[K]^{2}$ with probability at least $2/T$ as $K<<T$.
Next, in Lemma E.6 we bound the regret of Algorithm 8 for the case where $\lambda$ is unknown and small. More specifically, we are going to prove that Algorithm 8 after building the estimators $\hat{\mu}_{i}$ and $\hat{\mu}_{i,j}$ with the properties outlined above is going to call EXP3.P. The regret then follows from the results of Section 4.
@@
If $\lambda$ is in $[0,\tilde{\Theta}(1/T)]$, then Algorithm 8 incurs regret:
while being agnostic to $\lambda$.